# Onsager reciprocal relations for charge and spin transport in periodically driven systems

### v1

@@ I Introduction
Periodically driven systems have opened a new way for light-induced and light-controlled transport phenomena. In general, systems are periodically driven by a time-periodic field such as the pump field of light. Because of the time periodicity, the periodically driven systems are described by the Floquet Hamiltonian Floquet1 ; Floquet2 . Since this Hamiltonian can be tuned by varying the parameters of the pump field, various properties of the systems can be engineered without changing the materials. This is called Floquet engineering Floquet-review1 ; Floquet-review2 ; Floquet-review3 ; NA-Floquet-Mott1 . For example, the pump field of circularly polarized light (CPL) can be used to induce the anomalous Hall effect (AHE) Oka-PRB ; Mikami ; Light-AHE-exp1 ; Light-AHE-exp2 ; NA-FloquetSHE , in which the charge current perpendicular to the probe field is generated AHE-Hall ; Karplus-Luttinger ; AHE-review . Furthermore, the charge current generated in this AHE can be changed in magnitude and direction by varying the amplitude and helicity of CPL Oka-PRB ; Mikami ; Light-AHE-exp2 ; NA-FloquetSHE . Such optical control is also possible for the spin Hall effect (SHE) NA-FloquetSHE ; NA-TMD , in which the spin current, the flow of the spin angular momentum, is generated with the probe field perpendicular to it SHE-JETP ; SHE-Hirsch ; SHE-review .
@@
Despite various studies of transport phenomena in periodically driven systems, the Onsager reciprocal relations in these systems have been poorly understood. In general, the Onsager reciprocal relations connect two transport coefficients Onsager1 ; Onsager2 ; Kubo . Therefore, these relations provide general constraints on the symmetry of transport coefficients. Such constraints are useful for theoretical and experimental studies of transport phenomena. Moreover, the Onsager reciprocal relations are extremely important for spin transport. In general, it is much harder to detect the spin current than the charge current. Because of this, the spin current is usually detected indirectly in experiments. In fact, the observation of the inverse SHE can be regarded as the existence of the SHE InvSHE1 ; InvSHE2 because their transport coefficients are connected by the Onsager reciprocal relation; in the inverse SHE, the charge current perpendicular to the spin current is generated InvSHE1 ; InvSHE2 . Although there are some studies of the Onsager reciprocal relations in periodically driven systems Onsager-PD1 ; Onsager-PD2 ; Onsager-PD3 ; NA-Onsager , that relation for spin transport has been unexplored yet.
@@
In this paper, we theoretically study the Onsager reciprocal relations for charge and spin transport in periodically driven systems. The main results are summarized in Table 1. We consider the time-averaged charge and spin off-diagonal dc conductivities $\sigma_{yx}^{\textrm{C}}$ and $\sigma_{yx}^{\textrm{S}}$ to describe the charge and spin currents, respectively, perpendicular to the probe electric field in the linear-response regime for the nonequilibrium steady state of an electron system driven by the pump field of light. Figure 1 shows the set-up for their measurements. We begin with general arguments about the Onsager reciprocal relations for $\sigma_{yx}^{\textrm{C}}$ and $\sigma_{yx}^{\textrm{S}}$ in the systems driven by CPL, linearly polarized light (LPL), or bicircularly polarized light (BCPL). Here BCPL consists of a linear combination of the left- and right-handed CPL with different frequencies $\Omega$ and $\beta\Omega$ and a relative phase difference $\theta$ BCPL-NatPhoto ; BCPL-Oka ; NA-BCPL (see Fig. 2). We show that $\sigma_{yx}^{\textrm{C}}$ and $\sigma_{yx}^{\textrm{S}}$ satisfy the Onsager reciprocal relations in all the cases considered, although their main terms depend on the polarization of light. In the case with CPL or LPL, the main term of $\sigma_{yx}^{\textrm{C}}$ is given by the antisymmetric or symmetric part, respectively, whereas that of $\sigma_{yx}^{\textrm{S}}$ is given by the antisymmetric part. This suggests that the antisymmetric part of $\sigma_{yx}^{\textrm{S}}$ can be finite even with time-reversal symmetry, whereas that of $\sigma_{yx}^{\textrm{C}}$ is finite only without it. This difference arises from the difference between the time-reversal symmetries of the charge and spin currents. Meanwhile, in the case with BCPL, the Onsager reciprocal relations do not restrict $\sigma_{yx}^{\textrm{C}}$ and $\sigma_{yx}^{\textrm{S}}$ to either the antisymmetric or symmetric parts generally. This unusual property is due to the lack of a simple relation between the pump field of BCPL and its time-reversal counterpart. Then, we numerically test the Onsager reciprocal relations for $\sigma_{yx}^{\textrm{C}}$ and $\sigma_{yx}^{\textrm{S}}$ by applying the Floquet linear-response theory NA-FloquetSHE to a model of Sr2RuO4 driven by CPL, LPL, or BCPL with weak coupling to a heat bath. We demonstrate the validity of our arguments. Therefore, our results indicate that, even for periodically driven systems, the spin current can be detected by the inverse SHE. This is useful to develop and observe many spintronics phenomena in periodically driven systems. Then, our numerical calculations with BCPL show that the main term of $\sigma_{yx}^{\textrm{S}}$ is given by the antisymmetric part. Combining these results with the results with CPL or LPL, we conclude that $\sigma_{yx}^{\textrm{S}}$ can be regarded as the spin Hall conductivity in all the cases considered. Our numerical calculations with BCPL also show that the main term of $\sigma_{yx}^{\textrm{C}}$ depends on the magnitude, $\beta$, and $\theta$ of the pump field. More precisely, $\sigma_{yx}^{\textrm{C}}$ with BCPL for weak magnitude is dominated by the antisymmetric part, whereas that for moderately strong magnitude is almost vanishing in the cases of $(\beta,\theta)=(2,0)$, $(2,\frac{\pi}{2})$, $(2,\pi)$, $(3,0)$, and $(3,\pi)$ or dominated by the symmetric part in the cases of $(\beta,\theta)=(2,\frac{\pi}{4})$, $(2,\frac{3\pi}{4})$, $(3,\frac{\pi}{4})$, $(3,\frac{\pi}{2})$, and $(3,\frac{3\pi}{4})$. Therefore, $\sigma_{yx}^{\textrm{C}}$ cannot necessarily be regarded as the anomalous Hall conductivity even with broken time-reversal symmetry. Our results suggest that it is necessary to check the main terms of $\sigma_{yx}^{\textrm{C}}$ and $\sigma_{yx}^{\textrm{S}}$ in discussing the AHE and SHE, respectively.
@@
The remainder of this paper is organized as follows. In Sec. II, we argue the Onsager reciprocal relations in nondriven and periodically driven systems. After reviewing these relations in nondriven systems, we derive the Onsager reciprocal relations for $\sigma_{yx}^{\textrm{C}}$ and $\sigma_{yx}^{\textrm{S}}$ in the electron systems driven by CPL, LPL, or BCPL. In Sec. III, we introduce the model of periodically driven Sr2RuO4. In this model, we consider the heat bath as well as the system of Sr2RuO4 driven by the pump field of light and suppose that a nonequilibrium steady state is realized due to the damping induced by the coupling to the heat bath. In Sec. IV, we formulate $\sigma_{yx}^{\textrm{C}}$ and $\sigma_{yx}^{\textrm{S}}$ as well as their counterparts appearing in the Onsager reciprocal relations using the Floquet linear-response theory. We also comment on the applicability of this theory. In Sec. V, we show the numerical results of the Onsager reciprocal relations for $\sigma_{yx}^{\textrm{C}}$ and $\sigma_{yx}^{\textrm{S}}$ in Sr2RuO4 driven by CPL, LPL, or BCPL at $\beta=2$ and $3$. In Sec. VI, we discuss the origin of the characteristic $\theta$ dependences of $\sigma_{yx}^{\textrm{C}}$ and $\sigma_{yx}^{\textrm{S}}$ in the cases with BCPL, compare our results with other relevant studies, and remark on the experimental realization of our results. In Sec. VII, we make some concluding remarks on implications and outlooks.
@@
Table 1: The Onsager reciprocal relations for $\sigma_{yx}^{\textrm{C}}$ and $\sigma_{yx}^{\textrm{S}}$ in the nonequilibrium steady states of systems driven by CPL, LPL, and BCPL. $\sigma_{yx}^{\textrm{C}}$ and $\sigma_{yx}^{\textrm{S}}$ are the transport coefficients for the charge and spin currents along the $y$ axis, respectively, perpendicular to the probe electric field applied along the $x$ axis with the pump field $\mathbf{A}_{\textrm{pump}}(t)$; $\sigma_{xy}^{\textrm{C}}$ and $\tilde{\sigma}_{xy}^{\textrm{S}}$ are those for the charge currents along the $x$ axis perpendicular to the probe electric and spin fields, respectively, along the $y$ axis with $\mathbf{A}_{\textrm{pump}}(t)$; and $\bar{\sigma}_{xy}^{\textrm{C}}$ and $\bar{\tilde{\sigma}}_{xy}^{\textrm{S}}$ are those for the charge currents along the $x$ axis perpendicular to the probe electric and spin fields, respectively, along the $y$ axis with $\mathbf{A}_{\textrm{pump}}(-t)$. In the case with CPL or LPL, $\sigma_{yx}^{\textrm{C}}$ is restricted to the antisymmetric part $(\sigma_{yx}^{\textrm{C}}-\sigma_{xy}^{\textrm{C}})/2$ (corresponding to the anomalous Hall conductivity) or the symmetric part $(\sigma_{yx}^{\textrm{C}}+\sigma_{xy}^{\textrm{C}})/2$, respectively, whereas $\sigma_{yx}^{\textrm{S}}$ is restricted to the antisymmetric part $(\sigma_{yx}^{\textrm{S}}-\tilde{\sigma}_{xy}^{\textrm{S}})/2$ (corresponding to the spin Hall conductivity). (For the reason why we have defined the antisymmetric part of $\sigma_{yx}^{\textrm{S}}$ in this way, read the last paragraph of Sec. II B 1.) Meanwhile, in the case with BCPL, $\sigma_{yx}^{\textrm{C}}$ and $\sigma_{yx}^{\textrm{S}}$ are restricted to neither the antisymmetric nor symmetric parts generally.
@@
Figure 1: The setup for the pump-probe measurements of $\sigma_{yx}^{\textrm{C}}$ and $\sigma_{yx}^{\textrm{S}}$ in our periodically driven systems. $\sigma_{yx}^{\textrm{C}}$ and $\sigma_{yx}^{\textrm{S}}$ are the transport coefficients for describing the charge and spin currents, respectively, along the $y$ axis, which are perpendicular to the probe field applied along the $x$ axis in the nonequilibrium steady state with the pump field. The system is periodically driven Sr2RuO4 and is coupled to the heat bath. In Sr2RuO4, Ru4+ ions form the square lattice. This panel shows the case with the pump field of BCPL; the cases with those of CPL and LPL are also considered.
Figure 2: The trajectories of the pump field of BCPL at (a) $\beta=2$ and (b) $\beta=3$ and $\theta=0$, $\frac{\pi}{2}$, or $\pi$. The three or four loops with the same color are obtained per period of the pump field at $\beta=2$ or $3$, respectively.
II Onsager reciprocal relations
We begin with general argument about the Onsager reciprocal relations in nondriven or periodically driven systems. We argue these relations for the charge and spin off-diagonal dc conductivities in Secs. II A and II B, respectively. The results for periodically driven systems are summarized in Table 1. Although the Onsager reciprocal relations in nondriven systems may be well known, we will review them below to help the readers understand our general arguments in periodically driven systems. Since the essential symmetry in discussing the Onsager reciprocal relations is the symmetry against a time-reversal operation Onsager1 ; Onsager2 ; Kubo , we will not specify whether a certain vector is a vector or pseudo-vector.
@@ II Onsager reciprocal relations
Figure 3: The changes in (a) the charge current $\mathbf{J}_{\textrm{C}}$ and (b) the spin current $\mathbf{J}_{\textrm{S}}$ under the time-reversal operation. The up or down arrows represent the spin-up or spin-down electrons, respectively. Here $\mathbf{J}_{\textrm{C}}$ and $\mathbf{J}_{\textrm{S}}$ are defined as $\mathbf{J}_{\textrm{C}}=(-e)(\mathbf{J}_{\uparrow}+\mathbf{J}_{\downarrow})$ and $\mathbf{J}_{\textrm{S}}=(\hbar/2)(\mathbf{J}_{\uparrow}-\mathbf{J}_{\downarrow})$, where $\mathbf{J}_{\uparrow}$ and $\mathbf{J}_{\downarrow}$ are the currents of the spin-up and spin-down electrons, respectively. Under the time-reversal operation, $\mathbf{J}_{\uparrow}\rightarrow-\mathbf{J}_{\downarrow}$ and $\mathbf{J}_{\downarrow}\rightarrow-\mathbf{J}_{\uparrow}$, resulting in $\mathbf{J}_{\textrm{C}}\rightarrow-\mathbf{J}_{\textrm{C}}$ and $\mathbf{J}_{\textrm{S}}\rightarrow\mathbf{J}_{\textrm{S}}$.
Figure 4: The transport phenomena described by (a) $\sigma_{yx}^{\textrm{C}}=\sigma_{yx}^{\textrm{C}}[\mathbf{A}_{\textrm{pump}}(t)]$, (b) $\bar{\sigma}_{xy}^{\textrm{C}}=\sigma_{xy}^{\textrm{C}}[\mathbf{A}_{\textrm{pump}}(-t)]$, (c) $\sigma_{xy}^{\textrm{C}}=\sigma_{xy}^{\textrm{C}}[\mathbf{A}_{\textrm{pump}}(t)]$, (d) $\sigma_{yx}^{\textrm{S}}=\sigma_{yx}^{\textrm{S}}[\mathbf{A}_{\textrm{pump}}(t)]$, (e) $\bar{\tilde{\sigma}}_{xy}^{\textrm{S}}=\tilde{\sigma}_{xy}^{\textrm{S}}[\mathbf{A}_{\textrm{pump}}(-t)]$, and (f) $\tilde{\sigma}_{xy}^{\textrm{S}}=\tilde{\sigma}_{xy}^{\textrm{S}}[\mathbf{A}_{\textrm{pump}}(t)]$. The blue arrows represent the probe electric fields $E_{x}$ and $E_{y}$ and the probe spin field $E_{y}^{\textrm{S}}$. The orange arrows represent the charge currents $J_{\textrm{C}}^{y}$ and $J_{\textrm{C}}^{x}$ and the spin current $J_{\textrm{S}}^{y}$. The green arrows represent the pump fields $\mathbf{A}_{\textrm{pump}}(t)$ and $\mathbf{A}_{\textrm{pump}}(-t)$. The $x$ and $y$ axes are also drawn in these panels.
II.1 Charge off-diagonal dc conductivity

### v2

@@ 1 Introduction
Periodically driven systems have opened a new way for light-induced and light-controlled transport phenomena. In general, systems are periodically driven by a time-periodic field such as the pump field of light. Because of the time periodicity, the periodically driven systems are described by the Floquet Hamiltonian [1, 2]. Since this Hamiltonian can be tuned by varying the parameters of the pump field, various properties of the systems can be engineered without changing the materials. This is called Floquet engineering [3, 4, 5, 6]. For example, the pump field of circularly polarized light (CPL) can be used to induce the anomalous Hall effect (AHE) [7, 8, 9, 10, 11], in which the charge current perpendicular to the probe field is generated [12, 13, 14]. Furthermore, the charge current generated in this AHE can be changed in magnitude and direction by varying the amplitude and helicity of CPL [7, 8, 10, 11]. Such optical control is also possible for the spin Hall effect (SHE) [11, 15], in which the spin current, the flow of the spin angular momentum, is generated with the probe field perpendicular to it [16, 17, 18].
@@
Despite various studies of transport phenomena in periodically driven systems, the Onsager reciprocal relations in these systems have been poorly understood. In general, the Onsager reciprocal relations connect two transport coefficients [19, 20, 21]. Therefore, these relations provide general constraints on the symmetry of transport coefficients. Such constraints are useful for theoretical and experimental studies of transport phenomena. Moreover, the Onsager reciprocal relations are extremely important for spin transport. In general, it is much harder to detect the spin current than the charge current. Because of this, the spin current is usually detected indirectly in experiments. In fact, the observation of the inverse SHE can be regarded as the existence of the SHE [22, 23] because their transport coefficients are connected by the Onsager reciprocal relation; in the inverse SHE, the charge current perpendicular to the spin current is generated [22, 23]. Although there are some studies of the Onsager reciprocal relations in periodically driven systems [24, 25, 26, 27], that relation for spin transport has been unexplored yet.
@@
In this paper, we theoretically study the Onsager reciprocal relations for charge and spin transport in periodically driven systems. The main results are summarized in Table 1. We consider the time-averaged charge and spin off-diagonal dc conductivities $\sigma_{yx}^{\textrm{C}}$ and $\sigma_{yx}^{\textrm{S}}$ to describe the charge and spin currents, respectively, perpendicular to the probe electric field in the linear-response regime for the nonequilibrium steady state of an electron system driven by the pump field of light. Figure 1 shows the set-up for their measurements. We begin with general arguments about the Onsager reciprocal relations for $\sigma_{yx}^{\textrm{C}}$ and $\sigma_{yx}^{\textrm{S}}$ in the systems driven by CPL, linearly polarized light (LPL), or bicircularly polarized light (BCPL). Here BCPL consists of a linear combination of the left- and right-handed CPL with different frequencies $\Omega$ and $\beta\Omega$ and a relative phase difference $\theta$ [28, 29, 30] (see Fig. 2). We show that $\sigma_{yx}^{\textrm{C}}$ and $\sigma_{yx}^{\textrm{S}}$ satisfy the Onsager reciprocal relations in all the cases considered, although their main terms depend on the polarization of light. In the case with CPL or LPL, the main term of $\sigma_{yx}^{\textrm{C}}$ is given by the antisymmetric or symmetric part, respectively, whereas that of $\sigma_{yx}^{\textrm{S}}$ is given by the antisymmetric part. This suggests that the antisymmetric part of $\sigma_{yx}^{\textrm{S}}$ can be finite even with time-reversal symmetry, whereas that of $\sigma_{yx}^{\textrm{C}}$ is finite only without it. This difference arises from the difference between the time-reversal symmetries of the charge and spin currents. Meanwhile, in the case with BCPL, the Onsager reciprocal relations do not restrict $\sigma_{yx}^{\textrm{C}}$ and $\sigma_{yx}^{\textrm{S}}$ to either the antisymmetric or symmetric parts generally. This unusual property is due to the lack of a simple relation between the pump field of BCPL and its time-reversal counterpart. Then, we numerically test the Onsager reciprocal relations for $\sigma_{yx}^{\textrm{C}}$ and $\sigma_{yx}^{\textrm{S}}$ by applying the Floquet linear-response theory [11] to a model of Sr2RuO4 driven by CPL, LPL, or BCPL with weak coupling to a heat bath. We demonstrate the validity of our arguments. Therefore, our results indicate that, even for periodically driven systems, the spin current can be detected by the inverse SHE. This is useful to develop and observe many spintronics phenomena in periodically driven systems. Then, our numerical calculations with BCPL show that the main term of $\sigma_{yx}^{\textrm{S}}$ is given by the antisymmetric part. Combining these results with the results with CPL or LPL, we conclude that $\sigma_{yx}^{\textrm{S}}$ can be regarded as the spin Hall conductivity in all the cases considered. Our numerical calculations with BCPL also show that the main term of $\sigma_{yx}^{\textrm{C}}$ depends on the magnitude, $\beta$, and $\theta$ of the pump field. More precisely, $\sigma_{yx}^{\textrm{C}}$ with BCPL for weak magnitude is dominated by the antisymmetric part, whereas that for moderately strong magnitude is almost vanishing in the cases of $(\beta,\theta)=(2,0)$, $(2,\frac{\pi}{2})$, $(2,\pi)$, $(3,0)$, and $(3,\pi)$ or dominated by the symmetric part in the cases of $(\beta,\theta)=(2,\frac{\pi}{4})$, $(2,\frac{3\pi}{4})$, $(3,\frac{\pi}{4})$, $(3,\frac{\pi}{2})$, and $(3,\frac{3\pi}{4})$. Therefore, $\sigma_{yx}^{\textrm{C}}$ cannot necessarily be regarded as the anomalous Hall conductivity even with broken time-reversal symmetry. Our results suggest that it is necessary to check the main terms of $\sigma_{yx}^{\textrm{C}}$ and $\sigma_{yx}^{\textrm{S}}$ in discussing the AHE and SHE, respectively.
@@
The remainder of this paper is organized as follows. In Sect. 2, we argue the Onsager reciprocal relations in nondriven and periodically driven systems. After reviewing these relations in nondriven systems, we derive the Onsager reciprocal relations for $\sigma_{yx}^{\textrm{C}}$ and $\sigma_{yx}^{\textrm{S}}$ in the electron systems driven by CPL, LPL, or BCPL. In Sect. 3, we introduce the model of periodically driven Sr2RuO4. In this model, we consider the heat bath as well as the system of Sr2RuO4 driven by the pump field of light and suppose that a nonequilibrium steady state is realized due to the damping induced by the coupling to the heat bath. The reason why we choose Sr2RuO4 is twofold: its realistic model possesses the finite spin off-diagonal dc conductivity, which is more difficult to be realized than the finite charge off-diagonal dc conductivity; and its electronic structure is so simple that the Onsager reciprocal relations for these conductivities can be numerically studied even in the case with BCPL. In Sect. 4, we formulate $\sigma_{yx}^{\textrm{C}}$ and $\sigma_{yx}^{\textrm{S}}$ as well as their counterparts appearing in the Onsager reciprocal relations using the Floquet linear-response theory. We also comment on the applicability of this theory. In Sect. 5, we show the numerical results of the Onsager reciprocal relations for $\sigma_{yx}^{\textrm{C}}$ and $\sigma_{yx}^{\textrm{S}}$ in Sr2RuO4 driven by CPL, LPL, or BCPL at $\beta=2$ and $3$. In Sect. 6, we discuss the origin of the characteristic $\theta$ dependences of $\sigma_{yx}^{\textrm{C}}$ and $\sigma_{yx}^{\textrm{S}}$ in the cases with BCPL, compare our results with other relevant studies, and remark on the experimental realization of our results. In Sec. VII, we make some concluding remarks on implications and outlooks.
@@
Table 1: The Onsager reciprocal relations for $\sigma_{yx}^{\textrm{C}}$ and $\sigma_{yx}^{\textrm{S}}$ in the nonequilibrium steady states of systems driven by CPL, LPL, and BCPL. $\sigma_{yx}^{\textrm{C}}$ and $\sigma_{yx}^{\textrm{S}}$ are the transport coefficients for the charge and spin currents, respectively, along the $y$ axis perpendicular to the probe electric field applied along the $x$ axis with the pump field $\mathbf{A}_{\textrm{pump}}(t)$; $\sigma_{xy}^{\textrm{C}}$ and $\tilde{\sigma}_{xy}^{\textrm{S}}$ are those for the charge currents along the $x$ axis perpendicular to the probe electric and spin fields, respectively, along the $y$ axis with $\mathbf{A}_{\textrm{pump}}(t)$; and $\bar{\sigma}_{xy}^{\textrm{C}}$ and $\bar{\tilde{\sigma}}_{xy}^{\textrm{S}}$ are those for the charge currents along the $x$ axis perpendicular to the probe electric and spin fields, respectively, along the $y$ axis with $\mathbf{A}_{\textrm{pump}}(-t)$. In the case with CPL or LPL, $\sigma_{yx}^{\textrm{C}}$ is restricted to the antisymmetric part $(\sigma_{yx}^{\textrm{C}}-\sigma_{xy}^{\textrm{C}})/2$ (corresponding to the anomalous Hall conductivity) or the symmetric part $(\sigma_{yx}^{\textrm{C}}+\sigma_{xy}^{\textrm{C}})/2$, respectively, whereas $\sigma_{yx}^{\textrm{S}}$ is restricted to the antisymmetric part $(\sigma_{yx}^{\textrm{S}}-\tilde{\sigma}_{xy}^{\textrm{S}})/2$ (corresponding to the spin Hall conductivity). (For the reason why we have defined the antisymmetric part of $\sigma_{yx}^{\textrm{S}}$ in this way, read the last paragraph of Sect. 2.2.1.) Meanwhile, in the case with BCPL, $\sigma_{yx}^{\textrm{C}}$ and $\sigma_{yx}^{\textrm{S}}$ are restricted to neither the antisymmetric nor symmetric parts generally.
@@
Figure 1: (Color online) The setup for the pump-probe measurements of $\sigma_{yx}^{\textrm{C}}$ and $\sigma_{yx}^{\textrm{S}}$ in our periodically driven systems. $\sigma_{yx}^{\textrm{C}}$ and $\sigma_{yx}^{\textrm{S}}$ are the transport coefficients for describing the charge and spin currents, respectively, along the $y$ axis, which are perpendicular to the probe field applied along the $x$ axis in the nonequilibrium steady state with the pump field. The system is periodically driven Sr2RuO4 and is coupled to the heat bath. In Sr2RuO4, Ru4+ ions form the square lattice. This panel shows the case with the pump field of BCPL; the cases with those of CPL and LPL are also considered.
Figure 2: (Color online) The trajectories of the pump field of BCPL at (a) $\beta=2$ and (b) $\beta=3$ and $\theta=0$, $\frac{\pi}{2}$, or $\pi$. The three or four loops with the same color are obtained per period of the pump field at $\beta=2$ or $3$, respectively.
2 Onsager reciprocal relations
We begin with general argument about the Onsager reciprocal relations in nondriven or periodically driven systems. We argue these relations for the charge and spin off-diagonal dc conductivities in Sects. 2.1 and 2.2, respectively. The results for periodically driven systems are summarized in Table 1. Although the Onsager reciprocal relations in nondriven systems may be well known, we will review them below to help the readers understand our general arguments in periodically driven systems. Since the essential symmetry in discussing the Onsager reciprocal relations is the symmetry against a time-reversal operation [19, 20, 21], we will not specify whether a certain vector is a vector or pseudo-vector.
@@ 2 Onsager reciprocal relations
Figure 3: (Color online) The changes in (a) the charge current $\mathbf{J}_{\textrm{C}}$ and (b) the spin current $\mathbf{J}_{\textrm{S}}$ under the time-reversal operation. The up or down arrows represent the spin-up or spin-down electrons, respectively. Here $\mathbf{J}_{\textrm{C}}$ and $\mathbf{J}_{\textrm{S}}$ are defined as $\mathbf{J}_{\textrm{C}}=(-e)(\mathbf{J}_{\uparrow}+\mathbf{J}_{\downarrow})$ and $\mathbf{J}_{\textrm{S}}=(\hbar/2)(\mathbf{J}_{\uparrow}-\mathbf{J}_{\downarrow})$, where $\mathbf{J}_{\uparrow}$ and $\mathbf{J}_{\downarrow}$ are the currents of the spin-up and spin-down electrons, respectively. Under the time-reversal operation, $\mathbf{J}_{\uparrow}\rightarrow-\mathbf{J}_{\downarrow}$ and $\mathbf{J}_{\downarrow}\rightarrow-\mathbf{J}_{\uparrow}$, resulting in $\mathbf{J}_{\textrm{C}}\rightarrow-\mathbf{J}_{\textrm{C}}$ and $\mathbf{J}_{\textrm{S}}\rightarrow\mathbf{J}_{\textrm{S}}$.
Figure 4: (Color online) The transport phenomena described by (a) $\sigma_{yx}^{\textrm{C}}=\sigma_{yx}^{\textrm{C}}[\mathbf{A}_{\textrm{pump}}(t)]$, (b) $\bar{\sigma}_{xy}^{\textrm{C}}=\sigma_{xy}^{\textrm{C}}[\mathbf{A}_{\textrm{pump}}(-t)]$, (c) $\sigma_{xy}^{\textrm{C}}=\sigma_{xy}^{\textrm{C}}[\mathbf{A}_{\textrm{pump}}(t)]$, (d) $\sigma_{yx}^{\textrm{S}}=\sigma_{yx}^{\textrm{S}}[\mathbf{A}_{\textrm{pump}}(t)]$, (e) $\bar{\tilde{\sigma}}_{xy}^{\textrm{S}}=\tilde{\sigma}_{xy}^{\textrm{S}}[\mathbf{A}_{\textrm{pump}}(-t)]$, and (f) $\tilde{\sigma}_{xy}^{\textrm{S}}=\tilde{\sigma}_{xy}^{\textrm{S}}[\mathbf{A}_{\textrm{pump}}(t)]$. The blue arrows represent the probe electric fields $E_{x}$ and $E_{y}$ and the probe spin field $E_{y}^{\textrm{S}}$. Here $E_{y}^{\textrm{S}}$ is given by, for example, the gradient of either the Zeeman field or a spin-dependent chemical potential; it can induce a spin current as the probe electric field can induce a charge current. The orange arrows represent the charge currents $J_{\textrm{C}}^{y}$ and $J_{\textrm{C}}^{x}$ and the spin current $J_{\textrm{S}}^{y}$. The green arrows represent the pump fields $\mathbf{A}_{\textrm{pump}}(t)$ and $\mathbf{A}_{\textrm{pump}}(-t)$. The $x$ and $y$ axes are also drawn in these panels.
2.1 Charge off-diagonal dc conductivity